# DivScene: Benchmarking LVLMs for Object Navigation with Diverse Scenes and Objects

### v2

@@ 6.8 Case Study
In Figure 5, we provide an example of error analysis of our agent. The upper image shows the predicted path in the real scene, and the lower image shows the ground truth in the corresponding grid map. Here, the agent needs to find a soda can after walking through the whole game room. There are 46 actions in the shortest path. Instead of heading towards the goal, the agent just meanders in a limited area. This shows that our agent cannot finish the navigation with a long trajectory. One possible explanation could be the constraint of only providing the agent with recent historical information.
@@ 7 Conclusion
In this paper, we study a new task of building navigational agents capable of generalizing to a wide range of unseen objects in diverse scenes. For benchmarking, we collect a large-scale scene dataset, DivScene, featuring 81 different scene types. Then, we build the end-to-end agent, NatVLM, which is fine-tuned from an LVLM. The agent is trained through imitation learning on the shortest paths generated by a BFS-based planner. With the planner, we collect over 22K episodes to form the DivTraj dataset. In the training data, we also build CoT explanation traces to help the agent better grasp the underlying rationale of navigation. Extensive experiments show that our agent can navigate to diverse objects in diverse scenes significantly better than baselines. We also conduct various analyses to show the generalization ability of our agent. For future work, a promising direction is to expand the memory capacity of LVLMs to enable navigation over longer horizons.
@@
Our scene dataset DivScene is built upon the publicly available AI2THOR platform (Kolve et al., 2017) and the Holodeck framework (Yang et al., 2024). Then, we further extend our experiments on iTHOR (Weihs et al., 2021) and ProcTHOR (Deitke et al., 2022), both of which are open-source datasets.
Reproducibility
We provide the full implementation details of our NatVLM agent and baselines in Appendix D, including the learning rate, batch size, hard device, API access, etc. Meanwhile, we show the whole data postprocessing details in Sections B.3 and A.3.
Appendix A DivScene and DivTraj Details
@@ A.1 Scene Type Details and Attribute List
This section lists all scene types we collected by complementing the MIT scene dataset. In total, there are 81 scene types across five different categories, as shown in Table 6.
To collect diverse house descriptions, we add various attributes to a randomly sampled scene type. Here, we consider 12 different attributes as shown in Table 7. We ask GPT-4 to assign a value to each attribute and write a house description.
Table 6: The categories and scene types used in our DivScene dataset. In total, there are 5 categories and 81 scene types in our dataset.
@@ A.2 Textual House Description Prompt
We use in-context learning to prompt the GPT-4 to write textual house descriptions of 81 different scene types. Here, we show the concrete prompt we used in Table 8. We randomly sample 1-3 house attributes and ask GPT-4 to assign a value to them. Then, a house description is written based on the given scene type and attribute values. We use five exemplars in the in-context learning setting.
@@ A.3 Postprocessing of Textual House Description
After we collect textual house descriptions from GPT-4, we also introduce three filters to ensure their diversity and quality. First, we introduce a ROUGE-L filter. A new textual description is added only when its ROUGE-L similarity with any existing description is below 0.8, following previous works (Wang et al., 2022a; 2024a). Second, if we cannot find the values of given attributes in the first step of the output, we remove the example. Third, if we cannot find the house description in the second step in the output from GPT-4, we remove the example. The last two steps mean that the output does not follow the output format specified in the instructions and exemplars (see Table 8).
A.4 BFS-Based Planner
@@
The algorithm is shown in Algorithm 1, which is based on a priority queue. We design the BFS-based planner to pick the path with the fewest rotations to make it easier for LLMs to imitate. In detail, we add more costs when rotation changes since the agent needs one more rotation action before moving ahead, as shown at the 12th line in Algorithm 1.
@@
1:procedure bfs_search($reachable\_pos,start\_point,end\_point$, $start\_rotation$)
2:     Initialize priority queue $Q$, distance map $d\_map$, and parent map $p\_map$
3:     Enqueue $(0,start\_point,start\_rotation)$ to $Q$
@@
6:         if $current=end\_point$ then
7:              return ReconstructPath($p\_map$, $current$)
@@
[…]
10:              $neighbor\leftarrow$ GetNeighbor($current,r$)
@@
13:                  if $neighbor$ not visited or $newCost<$ $d\_map$[$neighbor$then
14:                       $d\_map[neighbor]=newCost$triangleright Update distance
15:                       $p\_map[neighbor]=current$triangleright Update parent
@@
22:end procedure
A.5 Action Derivation Algorithm
We show the action derivation algorithm in Algorithm 2. Between two adjacent positions in the shortest path, we add a MoveAhead action if the agent’s rotation remains unaltered. Otherwise, we first use RotateRight or RotateLeft to adjust the orientation and then add a MoveAhead action. After reaching the target object, we then rotate the agent so that the object is approximately centered in the agent’s egocentric view.
A.6 Data Diversity
To study what types of scenes are gathered under each category, we identify the category-type structure of houses in DivScene. We plot the top 10 most common scene types under each category in Figure 7. Then, we study the diversity of target objects in the episodes we sampled in DivTraj. We plot the top 15 most common scene types and their top 5 target object types in Figure 7. Overall, we see quite diverse scenes and target objects in our datasets.
@@ A.6 Data Diversity
1:procedure get_action_sequence($shortest\_path,start\_rotation,target\_object$)
2:     $agent\_rotation\leftarrow start\_rotation$
3:     Initialize empty $action\_list$
4:     $prior\_p\leftarrow shortest\_path[0]$
5:     for each $current\_p$ in $shortest\_path[1:]$ do
6:         $path\_rotation\leftarrow$ compute_path_rotation($current\_p$, $prior\_p$)
@@
8:              Append appropriate rotation action(s) to $action\_list$triangleright RotateRight or RotateLeft
@@
11:         Append “MoveAhead” to $action\_list$
@@
13:     $object\_rotation\leftarrow$ compute_object_rotation($target\_object,shortest\_path[-1]$)
@@
15:         Append final rotation action(s) to $action\_list$triangleright Adjust the view for the target object
@@
17:     Append “Done” to $action\_list$
18:     return $action\_list$
19:end procedure
@@ Appendix B NatVLM Instruction Data
In this section, we give the concrete prompts used in fine-tuning the LVLM, Idefics2.
@@ B.1 Instruction Template
We show the instruction template we used in the Table 9. There are four parts in the instruction template, including a brief task introduction, episode-specific information, the status of recent steps, and the prediction steps that need to be considered. We leave a lot of placeholders for the episode and step information.
@@ B.2 Response Template
Previous works usually collect explanation traces using GPT-4 (Mitra et al., 2023; Mukherjee et al., 2023; Ho et al., 2023). In contrast, we collect explanation traces with manually written templates. The templates and examples are shown in Tables 10 and 12. For RotateRight and RotateLeft, we identified the three most common scenarios for rotation based on heuristic rules. Then, we wrote the template for each of them. The first scenario is that the distance difference between the agent and the target object becomes zero in the agent’s rotation. Thus, the agent needs to navigate in the other direction. The second scenario involves the presence of obstacles in the agent’s current path, necessitating a rotation to navigate around them. The final scenario involves adjusting the agent’s rotation to center the target within its field of view, occurring at the end of the navigation process.
@@ B.3 Data Postprocessing
Directly using all actions in every trajectory to conduct imitation learning brings about an extremely imbalanced dataset. As shown in Table 11, 77.94% actions are MoveAhead. Then we downsampled MoveAhead actions in the instruction dataset. We only retain 25% of the MoveAhead actions resulting in a more balanced dataset.
@@
We show the final distribution of our dataset in Table 11 (w/ Postproc).
@@ C.2 Full Results of the Image Number Experiment
We provide the full results of the image number experiment of all metrics in Table 15(b).
C.3 Complementary Results of the Action Number Experiment
While we provide the results of the action number experiment with the default prompt in Table 3, we also provide the results with the difference-equation prompt in Table 13. The prompt is used as $\diamond$ w Diff-EQ” in the ablation study.
@@ C.3 Complementary Results of the Action Number Experiment
C.4 Complementary Results of the Few-Shot Experiment
While we provide the results of the few-shot learning with the default prompt in Table 4, we also provide the results with the difference-equation prompt in Table 14. The prompt is used as $\diamond$ w Diff-EQ” in the ablation study.
@@ C.4 Complementary Results of the Few-Shot Experiment
(a) Performance of the default prompt.
Number Valid Test All
SR SPL SEL SR SPL SEL SR SPL SEL
2 images 50.00 41.64 40.23 50.00 40.71 41.03 50.00 41.17 40.63
4 images 57.41 47.84 47.90 54.94 44.45 45.83 56.17 46.15 46.86
6 images 45.37 37.61 35.37 50.31 40.88 38.70 47.84 39.25 37.03
8 images 49.07 40.82 41.52 54.01 43.61 46.22 51.54 42.22 43.87
(b) Performance of the difference-equation prompt.
Number Valid Test All
SR SPL SEL SR SPL SEL SR SPL SEL
2 images 46.30 38.54 38.93 52.16 42.13 44.76 49.23 40.34 41.84
4 images 54.63 45.02 46.48 54.94 44.04 47.40 54.78 44.53 46.94
6 images 49.07 40.82 38.88 51.23 41.56 40.20 50.15 41.19 39.54
8 images 55.56 45.81 43.40 57.41 46.39 44.40 56.48 46.10 43.90
@@ Appendix E More House Examples
In this section, we present more houses built in our DivScene dataset in Figure 8.
@@
Appendix F Limitation
As we discussed in Section 6.8, our model fails to finish the navigation for long-horizon tasks. Instead of heading towards the goal, our NatVLM agent only meanders in a limited area near the start point. One possible explanation could be the constraint of only providing the agent with recent historical information. Thus, a promising direction for future work is to expand the memory capacity of LVLMs to enable navigation over longer horizons. For example, we can build a memory mechanism to help the LVLM-based agent maintain a long history of navigation. Meanwhile, we can better keep the original language and vision abilities of the LVLM used in our agent after fine-tuning, like the commonsense knowledge (Sap et al., 2019; Do et al., 2024; Wang et al., 2024b), dialogue (Yi et al., 2024; Choi et al., 2023), event understanding (Wang et al., 2022b; 2023b; Fang et al., 2024), etc (Cai et al., 2024a; Wang, 2022; Du et al., 2024).

### v3

@@ 6.8 Case Study
In Figure˜5, we provide an example of error analysis of NatVLM. The left image shows the predicted path in the real scene, and the right image shows the ground truth in the corresponding grid map. Here, the agent needs to find a soda can after walking through the whole room. There are 46 actions in the shortest path. Instead of heading towards the goal, the agent just meanders in a limited area. This shows that NatVLM cannot finish the navigation with a long trajectory and has limited exploration. We leave the development of models better at exploring environments for future work.
@@ 7 Conclusion
In this work, we first study the open-vocab object navigation and evaluate extensive LVLMs and LLMs. Meanwhile, we collect a large-scale scene dataset, DivScene, featuring 4,614 scenes. Over 22K episodes are sampled with the shortest paths by BFS. Our evaluation shows that current LVLMs and LLMs still fall short of the ability of open-vocab object navigation. Then, we also fine-tuned LVLMs with the shortest paths through imitation learning, where we also introduced CoT explanations. Experiments show that the navigation ability of LVLMs can be improved significantly. For future work, a promising direction is to enable LVLMs to navigate over longer horizons.
Limitation
While we conducted extensive experiments, our model still has some failure cases due to long navigation horizon or exploration, as we discussed in Section˜6.8. This might be because, due to the limited context window, we can only provide existing models with recent historical information. Thus, a promising direction for future work is to improve the memory (Du et al., 2025) or long-context (Team et al., 2025; Wang et al., 2025) capabilities of LVLMs to enable navigation over longer horizons.
@@
Our scene dataset DivScene is built upon the publicly available AI2THOR platform (Kolve et al., 2017) and the Holodeck framework (Yang et al., 2024). We further extend our experiments on iTHOR (Weihs et al., 2021) and ProcTHOR (Deitke et al., 2022), both of which are open-source datasets. We provide the complete details of the implementation of our NatVLM agent and baselines in Appendix˜D, including the learning rate, batch size, hard device, API access, etc. Meanwhile, we show the details of all the post-processing data in Sections˜B.3 and A.3.
Acknowledgements
The authors of this paper were supported by the ITSP Platform Research Project (ITS/189/23FP) from ITC of Hong Kong, SAR, China, and the AoE (AoE/E-601/24-N), the RIF (R6021-20) and the GRF (16205322) from RGC of Hong Kong, SAR, China. This work was done when Zhaowei Wang worked as an intern at Tencent AI Lab.
Appendix A DivScene and DivSceneep Details
@@ A.1 Scene Type Details and Attribute List
This section lists all scene types we collected by complementing the MIT scene dataset. In total, there are 81 scene types across five different categories, as shown in Table˜7.
To collect diverse house descriptions, we add various attributes to a randomly sampled scene type. Here, we consider 12 different attributes as shown in Table˜8. We ask GPT-4 to assign a value to each attribute and write a house description.
@@
Table 7: The categories and scene types used in our DivScene dataset. In total, there are five categories and 81 scene types in our dataset.
@@ A.2 Textual House Description Prompt
We use in-context learning to prompt the GPT-4 to write textual house descriptions of 81 different scene types. Here, we show the concrete prompt we used in Table˜9. We randomly sample 1-3 house attributes and ask GPT-4 to assign a value to them. Then, a house description is written based on the given scene type and attribute values. We use five exemplars in the in-context learning setting.
@@ A.3 Postprocessing of Textual House Description
After we collect textual house descriptions from GPT-4, we also introduce three filters to ensure their diversity and quality.
  • We first introduce a ROUGE-L filter. To encourage diversity, a new description is discarded when its ROUGE-L similarity (Lin, 2004) with any existing description is above 0.8, following previous works (Wang et al., 2022, 2024).
  • Second, if we cannot correctly parse the output in the first step and find the values of the given attributes, we remove the example.
  • Third, if we cannot find the house description in the second step in the output from GPT-4, we remove the example. The last two steps mean that the output does not follow the output format specified in the instructions and exemplars (see Table˜9).
A.4 Breadth-First Search for Shortest Path
@@
The algorithm is shown in Algorithm˜1, which is based on a priority queue. We design the BFS-based planner to pick the path with the fewest rotations to make it easier for LLMs to imitate. In detail, we add more costs when rotation changes since the agent needs one more rotation action before moving ahead, as shown at the 12th line in Algorithm˜1.
A.5 Action Derivation Algorithm
We show the action derivation algorithm in Algorithm˜2. Between two adjacent positions in the shortest path, we add a MoveAhead action if the agent’s rotation remains unaltered. Otherwise, we first use RotateRight or RotateLeft to adjust the orientation and then add a MoveAhead action. After reaching the target object, we then rotate the agent so that the object is approximately centered in the agent’s egocentric view.
A.6 Data Diversity
To study what types of scenes are gathered under each category, we identify the category-type structure of houses in DivScene. We plot the top 10 most common scene types under each category in Figure˜7. Then, we study the diversity of target objects in the episodes we sampled in DivSceneep. We plot the top 15 most common scene types and their top 5 target object types in Figure˜7. Overall, we see quite diverse scenes and target objects in our datasets.
@@ A.6 Data Diversity
1:Input: reachable_pos,start_point,end_point,start_rotationreachable\_pos,start\_point,end\_point,start\_rotation
2:Initialize priority queue $Q$, distance map d_mapd\_map, and parent map p_mapp\_map
3:Enqueue (0,start_point,start_rotation)(0,start\_point,start\_rotation) to $Q$
@@
6:  if current=end_pointcurrent=end\_point then
7:   return ReconstructPath(p_mapp\_map, currentcurrent)
@@
9:  for $r\in\{$North, East, South, West}\} do
10:   $neighbor\leftarrow$ GetNeighbor($current,r$) $\triangleright$ Four cardinal directions
@@
13:     if neighborneighbor not visited or newCost<d_map[neighbor]newCost<d\_map[neighbor] then
14:      d_map[neighbor]=newCostd\_map[neighbor]=newCost $\triangleright$ Update distance
15:      p_map[neighbor]=currentp\_map[neighbor]=current $\triangleright$ Update parent
@@
1:Input: shortest_path,start_rotation,target_objectshortest\_path,start\_rotation,target\_object
[…]
3:Initialize empty action_listaction\_list
[…]
5:for current_pcurrent\_p in shortest_path[1:]shortest\_path[1:] do
6:  $path\_rotation\leftarrow$ compute_path_rotation(current_pcurrent\_p, prior_pprior\_p)
@@
8:   Append appropriate rotation action(s) to action_listaction\_list $\triangleright$ RotateRight or RotateLeft
@@
11:  Append “MoveAhead” to action_listaction\_list
@@
[…]
@@
15:  Append final rotation action(s) to action_listaction\_list $\triangleright$ Adjust the view for the target object
@@
17:Append “Done” to action_listaction\_list
18:return action_listaction\_list
@@ Appendix B NatVLM Instruction Data
In this section, we give the concrete prompts used in fine-tuning NatVLM.
@@ B.1 Instruction Template
We show the instruction template we used in the Table˜10. There are four parts in the instruction template, including a brief task introduction, episode-specific information, the status of recent steps, and the prediction steps that need to be considered. We leave a lot of placeholders for the episode and step information.
@@ B.2 Response Template
Previous works usually collect explanation traces using GPT-4 (Mitra et al., 2023; Mukherjee et al., 2023; Ho et al., 2023). In contrast, we collect explanation traces with manually written templates. The templates and examples are shown in Tables˜11 and 13. For RotateRight and RotateLeft, we identified the three most common scenarios for rotation based on heuristic rules. Then, we wrote the template for each of them. The first scenario is that the distance difference between the agent and the target object becomes zero in the agent’s rotation. Thus, the agent needs to navigate in the other direction. The second scenario involves the presence of obstacles in the agent’s current path, necessitating a rotation to navigate around them. The final scenario involves adjusting the agent’s rotation to center the target within its field of view, occurring at the end of the navigation process.
@@ B.3 Data Postprocessing
Directly using all actions in every trajectory to conduct imitation learning brings about an extremely imbalanced dataset. As shown in Table˜12, 77.94% actions are MoveAhead. Then we downsampled MoveAhead actions in the instruction dataset. We only retain 25% of the MoveAhead actions resulting in a more balanced dataset.
@@
We show the final distribution of our dataset in Table˜12 (w/ Postproc).
@@ C.2 Full Results of the Image Number Experiment
We provide the full results of the image number experiment of all metrics in Table˜14. Besides the default prompt structure, we also conduct experiments with the difference-equation prompt, which is introduced as $\diamond$ w Diff-EQ” in the ablation study.
Number Valid Test All
SR SPL SEL SR SPL SEL SR SPL SEL
2 images 50.00 41.64 40.23 50.00 40.71 41.03 50.00 41.17 40.63
4 images 57.41 47.84 47.90 54.94 44.45 45.83 56.17 46.15 46.86
6 images 45.37 37.61 35.37 50.31 40.88 38.70 47.84 39.25 37.03
8 images 49.07 40.82 41.52 54.01 43.61 46.22 51.54 42.22 43.87
(a) Performance of the default prompt.
Number Valid Test All
SR SPL SEL SR SPL SEL SR SPL SEL
2 images 46.30 38.54 38.93 52.16 42.13 44.76 49.23 40.34 41.84
4 images 54.63 45.02 46.48 54.94 44.04 47.40 54.78 44.53 46.94
6 images 49.07 40.82 38.88 51.23 41.56 40.20 50.15 41.19 39.54
8 images 55.56 45.81 43.40 57.41 46.39 44.40 56.48 46.10 43.90
(b) Performance of the difference-equation prompt.
C.3 Full Results of the Action Number Experiment
We provide the full results of the action number experiment with the default prompt in Table˜15. We also provide the results with the difference-equation prompt in Table˜16. The prompt is used as $\diamond$ w Diff-EQ” in the ablation study.
Number Valid Test All
SR SPL SEL SR SPL SEL SR SPL SEL
4 steps 46.30 38.66 39.82 45.99 37.45 39.60 46.14 38.05 39.71
8 steps 57.41 47.84 47.90 54.94 44.45 45.83 56.17 46.15 46.86
12 steps 42.59 35.92 36.43 50.93 41.15 43.20 46.76 38.53 39.81
16 steps 51.85 43.01 42.72 53.40 43.03 44.08 52.62 43.02 43.40
Table 15: Hyperparameter investigation. Our agents receive various numbers of recent actions and positions within the default prompt.
@@ C.3 Full Results of the Action Number Experiment
C.4 Full Results of the Few-Shot Experiment
We provide the full results of the few-shot learning with the default prompt in Table˜17. We also provide the results with the difference-equation prompt in Table˜18. The prompt is used as $\diamond$ w Diff-EQ” in the ablation study.
% Data Valid Test Test Diff w GPT-4o
SR SPL SEL SR SPL SEL $\Delta_{SR}$ $\Delta_{SPL}$ $\Delta_{SEL}$
Table 17: Few-shot learning ability. We test our agent with different proportions of training data with the default prompt.
@@ C.4 Full Results of the Few-Shot Experiment
C.5 Zero-shot Transferring Evaluation
iTHOR and ProcTHOR both encompass four distinct scene types: bedrooms, living rooms, kitchens, and bathrooms. iTHOR has 30 rooms for each scene type, designed by professional 3D artists. Meanwhile, ProcTHOR is a procedural house-generation system that constructs 10,000 unique houses automatically. Similar to iTHOR, we sample 30 houses for each scene type from ProcTHOR. Four episodes are sampled in each house to evaluate our agent.
On the HM3D dataset, we test NatVLM on all houses. We compare our model with a lot of baselines on the HM3D dataset, including two lines of work: fine-tuned methods and zero-shot methods. As shown in Table˜19, the SPL scores of baselines are mostly lower than 30%. In contrast, the SPL score of NatVLM is 34.11%, 10 points higher than theirs.
Meanwhile, we want to emphasize that those LLM or VLM-based complicated navigation framework Ramrakhya et al. (2023); Majumdar et al. (2022); Zhou et al. (2023); Yokoyama et al. (2024a); Yin et al. (2024); Long et al. (2024) usually involve a lot more information than NatVLM. For example, our method only takes ego-centric RGB images as input, while those methods also get depth information, scene graphs, local policy navigation tools, extra VLMs (like GPT-4) besides backbone VLM, semantic segmentation, and panoramic field of views. This makes the comparison inherently unfair, as their methods rely on significantly more resources and additional modalities.
Meanwhile, we also do not include those complicated frameworks in other experiments due to the following reasoning: (1) They fall outside the evaluation scope of the current LVLMs; (2) They utilize different simulation platforms, such as Habitat Szot et al. (2021), rather than AI2THOR, which introduces significant compatibility issue with AI2THOR; (3) As abovementioned, they also introduce a lot of extra information beyond than ego-centric RGB, making the evaluation unfair.
Methods SPL SR
Fine-tuned
ProcTHOR Deitke et al. (2022) 31.8 54.4
PIRLNav Ramrakhya et al. (2023) 27.1 64.1
Zero-Shot
ProcTHOR (zero-shot) 7.7 13.2
ZSON Majumdar et al. (2022) 12.6 25.5
ESC Zhou et al. (2023) 22.3 39.2
VLFM Yokoyama et al. (2024a) 30.4 52.5
SG-Nav Yin et al. (2024) 24.8 53.9
InstructNav Long et al. (2024) 20.9 58.0
NatVLM (Our) 34.1 41.5
Table 19: Comparison of different models on HM3D, including fine-tuned and zero-shot methods.
@@ Appendix D Implementation Details
Here, we discuss the computational cost. For training, we use 8 $\times$ A100 GPUs to train our model. After adding the CoT traces, the training time increases from 10 hours to 17 hours, which is acceptable. Meanwhile, we use 1 A100 GPU to deploy the model. The generation time increases from 0.28s to 1.03s per query when we add the CoT traces. In practice, we can use more GPUs to speed up the inference. For example, RT-2 Brohan et al. (2023) uses a multi-TPU cloud service to achieve a frequency of 5Hz for a 5B model and 3Hz for a 55B model. The computational cost, both in training and inference, is reasonable in our experiments.
@@ Appendix E More House Examples
In this section, we present more houses built in our DivScene dataset in Figure˜9.
(a) a community center with versatile flooring
(b) a house from ProcTHOR
Figure 8: Comparing our houses with ProcTHOR.
@@
E.1 Comparison with ProcTHOR
In Figure˜8, we compare houses from our dataset and ProcTHOR with the same number of rooms (8 rooms). Obviously, our scene is more complex with more objects. Quantitively, our scene contains 466 objects, and the scene from ProcTHOR contains only 74 objects.